OpenStudy (anonymous):

cos(x+y)=0.17 and cosx=0.50, find siny

OpenStudy (dumbcow):

Use the angle sum formula for cosine \[\cos(x+y) = \cos(x) \cos(y) - \sin(x) \sin(y)\] Use pythagorean identity to find sin(x),cos(y) \[\sin^{2}(x) = 1-\cos^{2} (x)\] \[\cos^{2} (y) = 1-\sin^{2}(y)\]

OpenStudy (anonymous):

\[0.17=\frac{ cosy - \sqrt{3}siny }{ 2 }\] that is what i got after using sum formula for cosine.. i don't know what to do next

OpenStudy (dumbcow):

you need to replace cos(y) with an expression in terms of sin(y) since that is what you are trying to find look at pythagorean identity

OpenStudy (anonymous):

\[cosy=\sqrt{1-\sin ^{2}y}\] is that it?? i will substitute that?

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

\[0.34=\sqrt{1-\sin ^{2}y}-\sqrt{3}siny\] that's what i got.. and i don't know what to do next. will i square both sides of the equation?

OpenStudy (dumbcow):

you got it...i recommend moving the " sqrt3 siny" term to other side first

OpenStudy (dumbcow):

after you square both sides...combine like terms and set equal to zero then you will need to use quadratic formula

OpenStudy (anonymous):

ok.. i'll try to do that.. i'll post my answer afterwards and check it if its right.. thanks!! :))

OpenStudy (dumbcow):

ok yw

OpenStudy (gabylovesyou):

@dumbcow please help me!! :) http://openstudy.com/study#/updates/505dd105e4b02e1394100030

OpenStudy (anonymous):

i got a problem with my calculator :(( can you give me some site where i can best answer quadratic formula?? :((

OpenStudy (dumbcow):

wolframalpha.com will work as calculator http://www.wolframalpha.com/input/?i=4u^2+%2B.68sqrt%283%29u+%2B%28.34^2+-1%29+%3D+0 i replaced sin(y) with u

OpenStudy (anonymous):

its simple cosx=0.5 that eans 1/2 hence x=60 now cos(x+y)=0.17 cos(60+y)=0.17 60+y=arccos(0.17) y=arccos(0.17)-60 cosy=cos(arccos(0.17)-60)

OpenStudy (anonymous):

sriramkumar did it easier :O i got the answer right away.. thanks to you two.. :))